# Approximating Discrimination Within Models When Faced With Several Non-Binary Sensitive Attributes

### v1

@@ 3.3 Algorithmic effectiveness analysis of ApproxDist
Now we discuss the choice of hyper-parameters (i.e., $m_{1}$ and $m_{2}$) according to Eq. (15). In fact, Eq. (15) can be approximately written as $1-c\cdot n^{\frac{m_{1}}{k+1}}/m_{2}^{m_{1}}$. We can calculate the order of magnitude of $n^{\frac{m_{1}}{k+1}}/m_{2}^{m_{1}}$ by taking the logarithm:
$\small-\lambda\triangleq\lg\Big{(}n^{\frac{m_{1}}{k+1}}/m_{2}^{m_{1}}\Big{)}=m_{1}\bigg{(}\frac{\lg n}{k+1}-\lg m_{2}\bigg{)}\,.$ (22)
Therefore ApproxDist could reach an approximate solution with probability at least $(1-c\cdot 10^{-\lambda})$. In practice, we choose positive integers $m_{2}$ and $m_{1}$ such that $\lambda$ is reasonably large, ensuring that the algorithm will reach an approximate solution with high probability.
@@ 4 Empirical Results
In this section, we elaborate on our experiments to evaluate the effectiveness of the proposed HFM in Eq. (8) and ExtendDist in Algorithm 1, as well as ApproxDist in Algorithm 2. These experiments are conducted to explore the following research questions: RQ1. Compared with the state-of-the-art (SOTA) baseline fairness measures, does the proposed HFM capture the discriminative degree of one classifier effectively, and can it capture the discrimination level when facing several sensitive attributes with multiple values at the same time? Moreover, compared with the baselines, can HFM capture the discrimination level from both individual and group fairness aspects? RQ2. Can ApproxDist approximate the direct computation of distances in Eq. (4) and (5) precisely, and how efficient is ApproxDist compared with the direct computation of distances? And by extension, can ExtendDist approximate the direct computation of distances in Eq. (6) and (7) precisely, and how efficient is ExtendDist compared with the direct computation of distances? RQ3. Will the choice of hyper-parameters (that is, $m_{1}$ and $m_{2}$ in ApproxDist and ExtendDist) affect the approximation results, and if the answer is yes, how? Furthermore, we also discuss the limitations of the proposed approximation methods at the end of this section.
4.1 Experimental setups
@@
Five public datasets were adopted in the experiments: Ricci, [^3] https://rdrr.io/cran/Stat2Data/man/Ricci.html Credit, [^4] https://archive.ics.uci.edu/ml/datasets/statlog+(german+credit+data Income, [^5] https://archive.ics.uci.edu/ml/datasets/adult PPR, and PPVR. [^6] https://github.com/propublica/compas-analysis/, that is,
Propublica-Recidivism and Propublica-Violent-Recidivism datasets
Each of them has two sensitive attributes except Ricci, with more details provided in Table I below.
TABLE I: Dataset statistics. The column ‘#inst’ represents the number of instances; The $\text{joint}_{\text{both}}$ column represents that both two of the sensitive attribute values of one instance belong to the corresponding privileged group; The $\text{joint}_{\text{either}}$ column represents for this instance, at least one of its sensitive values belongs to the corresponding privileged group.
Dataset #inst #feature #member in the privileged group
raw processed 1st priv 2nd priv $\text{joint}_{\text{both}}$ $\text{joint}_{\text{either}}$
ricci 118 5 6 68 in race
credit 1000 21 58 690 in sex 851 in age 625 916
income 30162 14 98 25933 in race 20380 in sex 18038 28275
ppr 6167 11 401 4994 in sex 2100 in race 1620 5474
ppvr 4010 11 327 3173 in sex 1452 in race 1119 3506
@@
To evaluate the validity of HFM in capturing the discriminative degree of classifiers, we compare it with three commonly-used group fairness measures (that is, demographic parity (DP) [35, 36], equality of opportunity (EO) [37], and predictive quality parity (PQP) [38, 2]) [^7] Three commonly used group fairness measures of one classifier $f(\cdot)$ are evaluated as $\mathrm{DP}(f)$ $=\lvert\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!1]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!0]\rvert\,,$ (23a) $\mathrm{EO}(f)$ $=\lvert\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!1,\,y\!=\!1]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!0,\,y\!=\!1]\rvert,$ (23b) $\mathrm{PQP}(f)$ $=\lvert\mathbb{P}_{\mathcal{D}}[y\!=\!1|\mathop{\bm{a}}\!=\!1,\,f(\bm{x})\!=\!1]-\mathbb{P}_{\mathcal{D}}[y\!=\!1|\mathop{\bm{a}}\!=\!0,\,f(\bm{x})\!=\!1]\rvert,$ (23c) respectively, where $\bm{x}=(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})$, $y$, and $f(\bm{x})$ are respectively features, the true label, and the prediction of this classifier for one instance. Note that $\mathop{\bm{a}}=1$ and 00 respectively mean that the instance $\bm{x}$ belongs to the privileged group and marginalised groups. and one named discriminative risk (DR) [^8] The discriminative risk (DR) of this classifier is evaluated as $\scriptsize\mathrm{DR}(f)=\mathbb{E}_{\mathcal{D}}[\mathbb{I}(f(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})\neq f(\mathop{\breve{\bm{x}}},\mathop{\tilde{\bm{a}}}))]\,,$ (24) where $\mathop{\tilde{\bm{a}}}$ represents the disturbed sensitive attributes. DR reflects its bias degree from both individual- and group-fairness aspects. [32] that could reflect the bias level of ML models from both individual- and group-fairness aspects.
@@
TABLE II: Test evaluation performance of different fairness measures, where LightGBM is used as the learning algorithm. The column named ‘$\text{Att}_{\text{sen}}$’ denotes a corresponding sensitive attribute, and $\Delta$ denotes the performance difference between a metric and that after disturbing the data [32]. Note that $\mathbf{df}_{\text{prev}}=\nicefrac{{\mathbf{D}_{f}(S_{1},\bar{S}_{1})}}{{\mathbf{D}(S_{1},\bar{S}_{1})}}-1$ represents our previous work [34], and $\mathbf{df}=\log(\nicefrac{{\mathbf{D}_{f,\mathop{\bm{a}}}(S,a_{i})}}{{\mathbf{D}_{\mathop{\bm{a}}}(S,a_{i})}})$ and $\mathbf{df}^{\text{avg}}=\log(\nicefrac{{\mathbf{D}_{f,\mathop{\bm{a}}}^{\text{avg}}(S,a_{i})}}{{\mathbf{D}_{\mathop{\bm{a}}}^{\text{avg}}(S,a_{i})}})$ here represent HFM in this paper for each sensitive attribute.
@@
TABLE III: Test evaluation performance of different fairness measures, where LightGBM is used as the learning algorithm. The notation $\Delta$ denotes the performance difference between a metric and that after disturbing the data and DR works for one sensitive attribute with multiple values [32]. Here we use $\text{DR}_{avg}=\frac{1}{n_{a}}\sum_{i=1}^{n_{a}}\text{DR}_{i}$ to reflect the bias level on the whole dataset.
@@
Figure 1: Correlation heatmap between normal evaluation metric and fairness, for one single sensitive attribute. The used notations refer to those in Table III.
Figure 2: Correlation heatmap between normal evaluation metric and fairness measure, for all sensitive attributes within the dataset. The notations used here refer to those in Table III.
Figure 3: Plots of best test-set fairness-performance trade-offs per fairness metric [39] (the smaller the better). (a) Plot of fairness-accuracy trade-off for one single sensitive attribute; (b) Plot of fairness-accuracy trade-off for all sensitive attributes; (c–d) Plots of fairness-$\mathrm{f}_{1}$ score trade-off for one sensitive attribute and for all sensitive attributes, respectively. Note that the notations in (a) and (c) refer to those in Table III, and that in (b) and (d) refer to those in Table III.
@@
Figure 5: Effect of hyper-parameters $m_{1}$ and $m_{2}$ in ApproxDist and ExtendDist. (a–b) The effect of $m_{2}$ on the maximal distance value; (c–d) The effect of $m_{1}$ on the maximal distance values. (e–f) The effect of $m_{2}$ on the average distance value; (g–h) The effect of $m_{1}$ on the average distance value. (i–j) The effect of $m_{2}$ on the time cost; (k–l) The effect of $m_{1}$ on the time cost, where $m_{2}$ is set to $\lceil 2\lg(n)\rceil$ in terms of $n$—the size of the corresponding dataset.
@@
Figure 6: Effect of hyper-parameters $m_{1}$ and $m_{2}$ in our previous work [34] and ApproxDist here, where only binary values are considered for sensitive attributes. (a–b) The effect of $m_{2}$ on the (maximal) distance value; (c–d) The effect of $m_{2}$ on the time cost. (e–f) The effect of $m_{1}$ on the (maximal) distance value; (g–h) The effect of $m_{1}$ on the time cost, where $m_{2}$ is fixed to $\lceil 2\lg(n)\rceil$ in terms of $n$—the size of the corresponding dataset.

### v2

@@ 3.3 Algorithmic effectiveness analysis of ApproxDist
Now we discuss the choice of hyperparameters (i.e. $m_{1}$ and $m_{2}$) according to Eq. (13). In fact, Eq. (13) can be approximately written as $1-c\cdot n^{\frac{m_{1}}{k+1}}/m_{2}^{m_{1}}$. We can calculate the order of magnitude of $n^{\frac{m_{1}}{k+1}}/m_{2}^{m_{1}}$ by taking the logarithm:
$\small-\lambda\triangleq\lg\Big{(}n^{\frac{m_{1}}{k+1}}/m_{2}^{m_{1}}\Big{)}=m_{1}\Big{(}\tfrac{\lg n}{k+1}-\lg m_{2}\Big{)}\,.$ (20)
Therefore, ApproxDist could reach an approximate solution with probability at least $(1-c\cdot 10^{-\lambda})$. In practice, we choose positive integers $m_{2}$ and $m_{1}$ such that $\lambda$ is reasonably large, ensuring that the algorithm will reach an approximate solution with high probability.
@@ 4 Empirical Results
In this section, we elaborate on our experiments to evaluate the effectiveness of the proposed HFM in Eq. (8) and ExtendDist in Algorithm 1, as well as ApproxDist in Algorithm 2. These experiments are conducted to explore the following research questions: RQ1. Compared with the state-of-the-art (SOTA) baseline fairness measures, does the proposed HFM capture the discriminative degree of one classifier effectively, and can it capture the discrimination level when facing several SAs with multiple values at the same time? Moreover, compared with the baselines, can HFM capture the discrimination level from both individual and group fairness aspects? RQ2. Can ApproxDist approximate the direct computation of distances in Eq. (4) and (5) precisely, and how efficient is ApproxDist compared with the direct computation of distances? And by extension, can ExtendDist approximate the direct computation of distances in Eq. (6) and (7) precisely, and how efficient is ExtendDist compared with the direct computation of distances? RQ3. Will the choice of hyperparameters (that is, $m_{1}$ and $m_{2}$ in ApproxDist and ExtendDist) affect the approximation results, and if the answer is yes, how? Furthermore, we also discuss the possibility of applying HFM to represented features in neural networks and the limitations of the proposed approximation methods at the end of this section.
4.1 Experimental setup
@@
Five public datasets were adopted in the experiments; Each of them has two SAs except Ricci, with more details provided in Table II below.
TABLE II: Data statistics. The column ‘#inst’ represents the number of instances; ‘#feat’ represents the number of features (incl. one or two SAs) excluding classification labels, and ‘prep’ is the number of features after preprocessing. For each SA, ‘#val’ means the number of its values; ‘#in-priv’ is the number of members in the privileged group accordingly.
Dataset #inst #feat 1st sensitive attribute 2nd sensitive attribute
raw prep #val #in-priv #val #in-priv
ricci [37] 118 5 6 race 3 68
credit [38] 1000 20 58 sex 2 690 age 2 851
income [39] 30162 13 98 race 5 25933 sex 2 20380
ppr [40] 6167 10 401 sex 2 4994 race 6 2100
ppvr [40] 4010 10 327 sex 2 3173 race 6 1452
@@
To evaluate the validity of HFM in capturing the discriminative degree of classifiers, we compare it with three commonly-used group fairness measures (that is, demographic parity (DP) [41, 42], equality of opportunity (EO) [43], and predictive quality parity (PQP) [44, 2]) [^6] Three commonly used group fairness measures of one classifier $f(\cdot)$ are evaluated on only one bi-valued SA as $\mathrm{DP}(f)$ $=\lvert\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!1]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!0]\rvert\,,$ (21a) $\mathrm{EO}(f)$ $=\lvert\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!1,\,y\!=\!1]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1|\mathop{\bm{a}}\!=\!0,\,y\!=\!1]\rvert,$ (21b) $\mathrm{PQP}(f)$ $=\lvert\mathbb{P}_{\mathcal{D}}[y\!=\!1|\mathop{\bm{a}}\!=\!1,\,f(\bm{x})\!=\!1]-\mathbb{P}_{\mathcal{D}}[y\!=\!1|\mathop{\bm{a}}\!=\!0,\,f(\bm{x})\!=\!1]\rvert,$ (21c) respectively, where $\bm{x}=(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})$, $y$, and $f(\bm{x})$ are respectively the features, the true label, and the prediction of this classifier for one instance. Note that $\mathop{\bm{a}}=1$ and 00 respectively mean that the instance $\bm{x}$ belongs to the privileged group and the marginalised groups. , one extended fairness measure (called statistical parity (SP), inspired by [32, 33]) [^7] The original statistical parity (SP) in [32, 33] is neither a quantitative measure nor applied directly to multiple SAs. It is more like, for at least one (bi- or multi-valued) SA, that is, $\mathop{\bm{a}}\!=\![a_{1},...,a_{n_{a}}]^{\mathsf{T}}(n_{a}\!\geqslant\!1)$, $\mathrm{SP}^{\text{pre,max}}(f)$ $=\textstyle\max_{k\in\mathcal{A}^{\text{pre}}}\{|\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1\mid\mathop{\bm{a}}\!=\!k]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1]|\}\,,$ (22a) $\mathrm{SP}^{\text{pre,avg}}(f)$ $=\textstyle\sum_{k\in\mathcal{A}^{\text{pre}}}|\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1\mid\mathop{\bm{a}}\!=\!k]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1]|\,,$ (22b) where $\mathop{\bm{a}}\in\mathcal{A}^{\text{pre}}$ and $\mathcal{A}^{\text{pre}}$ is a finite set with $\prod_{1\leqslant i\leqslant n_{a}}|n_{a_{i}}|$ elements. It is not exactly the same but consistent with [32, 33]. For comparison, we made slight modifications following its original idea. For one multi-valued SA (that is, $a_{i}\!\in\!\{1,2,...,n_{a_{i}}\}$), two versions of the extended SP are $\mathrm{SP}_{i}^{\text{max}}(f)$ $=\textstyle\max_{1\leqslant j\leqslant n_{a_{i}}}\{|\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1\!\mid\!a_{i}\!=\!j]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1]|\}\,,$ (23a) $\mathrm{SP}_{i}^{\text{avg}}(f)$ $=\textstyle\sum_{1\leqslant j\leqslant n_{a_{i}}}|\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1\!\mid\!a_{i}\!=\!j]-\mathbb{P}_{\mathcal{D}}[f(\bm{x})\!=\!1]|\,,$ (23b) respectively. For several multi-valued SAs (that is, $\mathop{\bm{a}}\!=\![a_{1},...,a_{n_{a}}]^{\mathsf{T}}$), two versions of the extended SP are $\mathrm{SP}^{\text{max}}(f)$ $=\textstyle\max_{1\leqslant i\leqslant n_{a}}\mathrm{SP}_{i}^{\text{max}}(f)\,,$ (24a) $\mathrm{SP}^{\text{avg}}(f)$ $=\textstyle\frac{1}{n_{a}}\sum_{1\leqslant i\leqslant n_{a}}\mathrm{SP}_{i}^{\text{avg}}(f)\,,$ (24b) respectively. Neither (23) nor (24) is exactly the same as the original SP., and one named discriminative risk (DR) [^8] The discriminative risk (DR) of this classifier is evaluated as $\scriptsize\mathrm{DR}(f)=\mathbb{E}_{\mathcal{D}}[\mathbb{I}(f(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})\neq f(\mathop{\breve{\bm{x}}},\mathop{\tilde{\bm{a}}}))]\,,$ (25) where $\mathop{\tilde{\bm{a}}}$ represents the perturbed SAs. DR reflects its bias level from both individual- and group-fairness aspects. [34] that could reflect the bias level of ML models from both individual- and group-fairness aspects. Note that only the extended SP and DR can work for several multi-valued SAs.
@@
TABLE III: Test evaluation performance of different fairness measures, where LightGBM is used as the learning algorithm. The column named ‘$\text{Att}_{\text{sen}}$’ denotes a corresponding SA, and the notation $\Delta$(performance) denotes the performance difference between a metric and that after perturbing the data [34]. Note that $\mathbf{df}_{\text{prev}}=\nicefrac{{\mathbf{D}_{f}(S_{1},\bar{S}_{1})}}{{\mathbf{D}(S_{1},\bar{S}_{1})}}-1$ represents our previous work [36], and $\mathbf{df}=\log(\nicefrac{{\mathbf{D}_{f,\mathop{\bm{a}}}(S,a_{i})}}{{\mathbf{D}_{\mathop{\bm{a}}}(S,a_{i})}})$ and $\mathbf{df}^{\text{avg}}=\log(\nicefrac{{\mathbf{D}_{f,\mathop{\bm{a}}}^{\text{avg}}(S,a_{i})}}{{\mathbf{D}_{\mathop{\bm{a}}}^{\text{avg}}(S,a_{i})}})$ here represent HFM in this paper for each SA.
@@
TABLE IV: Test evaluation performance of different fairness measures, where LightGBM is used as the learning algorithm. The notation $\Delta$ denotes the performance difference between a metric and that after perturbing the data. Here $\text{DR}_{i}$ works for one SA with multiple values [34], and we use $\text{DR}_{\text{avg}}=\frac{1}{n_{a}}\sum_{i=1}^{n_{a}}\text{DR}_{i}$ to reflect the bias level on the whole dataset.
@@
Figure 1: Correlation heatmap between normal evaluation metric and fairness, for one single SA. The used notations refer to those in Table IV.
Figure 2: Correlation heatmap between normal evaluation metric and fairness measure, for all SAs within the dataset. The notations used here refer to those in Table IV.
Figure 3: Comparison of HFM with the extended SP. (a–b) Correlation heatmap between normal evaluation metric and fairness measure, for all SAs within the dataset. Note that in the computation of $\mathbf{df}_{\text{prev}}$ here, multi-valued SAs are handled as bi-valued cases, equivalent to $\mathbf{df}_{\text{prev\;bin-val}}$ in Figure 1; $\hat{\mathbf{df}}_{\text{prev}},\hat{\mathbf{df}}$, and $\hat{\mathbf{df}}^{\text{avg}}$ indicate that the distances are obtained using approximation algorithms. (c–d) Plots of best test-set fairness-accuracy trade-offs per fairness metric [45] (the smaller the better).
Figure 4: Plots of best test-set fairness-performance trade-offs per fairness metric [45] (the smaller the better). (a) Plot of fairness-accuracy trade-off for one single SA; (b) Plot of fairness-accuracy trade-off for all SAs; (c–d) Plots of fairness-$\mathrm{f}_{1}$ score trade-off for one SA and for all SAs, respectively. Note that the notations in (a) and (c) refer to those in Table IV, and that in (b) and (d) refer to those in Table IV.
@@
Figure 6: Effect of hyperparameters $m_{1}$ and $m_{2}$ in ApproxDist and ExtendDist. (a–b) The effect of $m_{2}$ on the maximal distance value; (c–d) The effect of $m_{1}$ on the maximal distance values. (e–f) The effect of $m_{2}$ on the average distance value; (g–h) The effect of $m_{1}$ on the average distance value. (i–j) The effect of $m_{2}$ on the time cost; (k–l) The effect of $m_{1}$ on the time cost, where $m_{2}$ is set to $\lceil 2\lg(n)\rceil$ in terms of $n$—the size of the corresponding dataset.
@@
Figure 7: Effect of hyperparameters $m_{1}$ and $m_{2}$ in our previous work [36] and ApproxDist here, where only binary values are considered for SAs. (a–b) The effect of $m_{2}$ on the (maximal) distance value; (c–d) The effect of $m_{2}$ on the time cost. (e–f) The effect of $m_{1}$ on the (maximal) distance value; (g–h) The effect of $m_{1}$ on the time cost, where $m_{2}$ is fixed to $\lceil 2\lg(n)\rceil$ in terms of $n$—the size of the corresponding dataset.
Implementation details
We mainly use bagging, AdaBoost, LightGBM [46], FairGBM [45], and AdaFair [47] as learning algorithms, where FairGBM and AdaFair are two fairness-aware ensemble-based methods. Plus, certain kinds of classifiers are used in Section 4.2—including decision trees (DT), naive Bayesian (NB) classifiers, $k$-nearest neighbours (KNN) classifiers, Logistic Regression (LR), support vector machines (SVM), linear SVMs (linSVM), and multilayer perceptrons (MLP)—so that we have a larger learner pool to choose from based on different fairness-relevant rules. Standard 5-fold cross-validation is used in these experiments; in other words, in each iteration, the entire dataset is divided into two parts, with 80% as the training set and 20% as the test set. Also, features of datasets are scaled in preprocessing to lie between 0 and 1. Except for the experiments for RQ3, we set the hyperparameters $m_{1}=25$ and $m_{2}=\lceil 2\lg(n)\rceil$ in all other experiments. The value $25$ of $m_{1}$ is randomly chosen, just as a value that is not too small to get an awful approximation meanwhile not too large to cost unnecessarily much more time. We also discuss, in RQ3 and Section 4.4, whether the choice of $m_{1}$ and $m_{2}$ will affect the approximation results.
4.2 Comparison between HFM and baseline fairness measures
The aim of this experiment is to evaluate the effectiveness of the proposed HFM compared with baseline fairness measures. As ground-truth discriminative levels of classifiers remain unknown and it is hard to directly compare different methods from that perspective, we compare the correlation (referring to the Pearson correlation coefficient) between the performance difference and different fairness measures. The empirical results are reported in Fig. 15 and Tables IVIV.
4.2.1 Comparison with three group fairness measures
For one single SA, we can see from Fig. 1 that: even $\mathbf{df}^{\text{avg}}$ only describes the extra bias, its correlation with $\Delta($performance)))) is still close to that of DR (and sometimes DP), which means HFM can capture the bias within classifiers indeed and that HFM captures it more finely than our previous work [36]. Moreover, $\mathbf{df}^{\text{avg}}$ shows higher correlation with $\Delta($performance)))) than $\mathbf{df}$ in most cases, which means $\mathbf{df}^{\text{avg}}$ may capture the extra bias level of classifiers better than $\mathbf{df}$ in practice.
As for multiple SAs, we can see from Fig. 2 that $\mathbf{df}^{\text{avg}}$ shows higher correlation with $\Delta($performance)))) than $\mathbf{df}$ in most cases, which is similar to our observation in Fig. 1. Note that the original DR [34] calculates all SAs with binary or multiple values as a whole, and for comparison with HFM, we calculate here $\text{DR}_{i}$ for each SA and $\text{DR}_{\text{avg}}=\frac{1}{n_{a}}\sum_{i=1}^{n_{a}}\text{DR}_{i}$, analogously to $\mathbf{df}^{\text{avg}}$. Besides, we observe that the correlation between $\mathbf{df}^{\text{avg}}$ and $\Delta\text{Accuracy}$ (resp. $\Delta\mathrm{f}_{1}$ score, $\Delta\text{Specificity}$) achieves half of that of DR, and $\mathbf{df}^{\text{avg}}$ even outperforms DR concerning $\Delta\text{Recall}$. Given that HFM only captures the extra bias introduced by classifiers, we believe that at least $\mathbf{df}^{\text{avg}}$ could capture quite a part of the bias within.
Furthermore, we report plots of fairness-performance trade-offs per fairness measure in Fig. 5. We can see that: 1) for one single SA, HFM (i.e. $\mathbf{df}$ and $\mathbf{df}^{\text{avg}}$) achieves the best result in Fig. 5LABEL:sub@subfig:tradeoff,a and 5LABEL:sub@subfig:tradeoff,c; and 2) for all SAs on one dataset, $\mathbf{df}$ and $\mathbf{df}^{\text{avg}}$ perform closely and both outperform $\text{DR}_{\text{avg}}$ in Fig. 5LABEL:sub@subfig:tradeoff,b and 5LABEL:sub@subfig:tradeoff,d. This observation demonstrates the effectiveness of HFM from another perspective, in other words, HFM could work well if fairness-performance trade-offs need to be considered.
4.2.2 Comparison with the extended SP
In Fig. 3LABEL:sub@subfig:extSP,b, $\mathbf{df}^{\text{avg}}$ and $\hat{\mathbf{df}}^{\text{avg}}$ achieve higher correlation with $\Delta$Accuracy, $\Delta$Precision, and $\Delta$Specificity, compared with the extended $\mathrm{SP}^{\text{max}}$. Although they do not have as high a correlation as the extended $\mathrm{SP}^{\text{avg}}$, their correlations are close to that of the extended $\mathrm{SP}^{\text{avg}}$, meaning that HFM can capture the discrimination at least partially, consistent with the extra discrimination that it is supposed to capture. Besides, Fig. 3LABEL:sub@subfig:extSP,c and 3LABEL:sub@subfig:extSP,d show that HFM (i.e. both $\mathbf{df}$ and $\mathbf{df}^{\text{avg}}$, as well as their approximated values) achieves better fairness-accuracy trade-offs than the extended SP, which is beneficial.
In this subsection, we evaluate the performance of the proposed ApproxDist and ExtendDist compared with the precise distance that is directly calculated by definitions. To verify whether they could achieve the true distance between sets in a precise and timely manner, we employ scatter plots to compare their values and time cost, presented in Fig. 5. Note that $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ and $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ are computed together in ApproxDist at one time, and so are $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ and $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ in ExtendDist. Notice that the previous ApproxDist [36] is included for comparison to its current version in scenarios of binary values. For comparison, we also use one more baseline named EarlyBreak [48], which is an efficient algorithm for calculating the exact Hausdorff distance (HD), that is, the distance of sets in this paper.
4.3.1 Validity of ApproxDist
As we can see from Figures 5LABEL:sub@subfig:approx,3 and 5LABEL:sub@subfig:approx,4, the approximated values of maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S,a_{i})$ are highly correlated with their corresponding precise values. Besides, their linear fit line and the identity line (that is, $f(x)=x$) are near and almost parallel, which means the approximated values are pretty close to their precise value. Similar observations are concluded for the average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S,a_{i})$ shown in Figures 5LABEL:sub@subfig:approx,7 and 5LABEL:sub@subfig:approx,8. As for the execution time of approximation and direct computation in Figures 5LABEL:sub@subfig:approx,11 and 5LABEL:sub@subfig:approx,12, ApproxDist may take a bit longer time in scenarios of multi-value cases than that of binary values, while all of them could achieve a shorter time than precise values when the execution of direct computation is costly.
4.3.2 Validity of ExtendDist
As we can see from Figures 5LABEL:sub@subfig:approx,1 and 5LABEL:sub@subfig:approx,2, the approximated values of maximal distance $\mathbf{D}_{\mathbf{a}}^{\text{}}(S)$ are highly correlated with their corresponding precise values. Besides, their linear fit line and the identity line are near and almost parallel, which means the approximated values are pretty close to their precise value. Similar observations are concluded for the average distance $\mathbf{D}_{\mathbf{a}}^{\text{avg}}(S)$ shown in Figures 5LABEL:sub@subfig:approx,5 and 5LABEL:sub@subfig:approx,6. As for the execution time of approximation and direct computation in Figures 5LABEL:sub@subfig:approx,9 and 5LABEL:sub@subfig:approx,10, ExtendDist would obtain a bigger advantage when computing precise values is expensive, while on the opposite, we do not need ExtendDist that much and can directly calculate them instead.
4.3.3 Comparison with EarlyBreak [48]
We use EarlyBreak [48] to compare with ApproxDist, in order to evaluate whether ApproxDist can efficiently approximate the exact values of distances. The reason why we do not compare it with ExtendDist is that EarlyBreak and the original exact Hausdorff distance are only for single-SA cases. The empirical results are provided in Fig. 9. We can observe that: (1) Both EarlyBreak and ApproxDist can achieve close computational or approximated values to the exact values of distances, with shorter time; (2) ApproxDist usually requires less time cost than EarlyBreak, shown in Fig. 9LABEL:sub@subfig:efficient,b. Thus, ApproxDist can at least serve as an acceptable solution even though $\mathcal{O}(n\log n)$ may still pose an absolute computational cost challenge as $n$ increases.